[Brahms-l] Shift report 20050520 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri May 20 2005 - 16:02:46 EDT
Shifter leader was HI
Shifters were EB


1) approx 3 hrs beam time;  "machine development" most of afternoon.
2) see E-log.
3) terminal for D1-D5 control "hanging"; reboot, log on to acnsun65.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri May 20 16:03:02 2005

This archive was generated by hypermail 2.1.8 : Fri May 20 2005 - 16:03:07 EDT